# Patronus: Bringing Transparency to Diffusion Models with Prototypes

### v1

@@ 4 Experiments and Results
In our experiments, we first assess the semantic meaning of learned prototypes (Sec. 4.1), by evaluating their ability to capture and reconstruct meaningful visual patterns through interpolation, extrapolation (prototype visualization), and manipulation tasks. We then analyze prototype quality and generation performance (Sec. 4.24.3). Finally, we explore how Patronus aids in diagnosing diffusion models by identifying potentially unwanted correlations via prototype similarity scores (Sec. 4.4).
We use five datasets: Quantitative evaluation on FMNIST [36], Cifar10 [15], FFHQ [13]; qualitative analysis on CheXpert [12]; and in-depth quantitative and qualitative experiments on CelebA [21]. We use 100 prototypes for all datasets except FMNIST, which has 30. Prototypes are encoded with a shape of (1,1,128). See Appendix for details.
@@
Capturing high level semantic features.
To check that the learned prototypes are semantically meaningful, we extract the prototype activation vector $s=Enc(x_{0})$ and sample random noise $x_{T}\sim\mathcal{N}(0,\mathbf{I})$ to generate new images $\hat{x}(s,x_{T})$. Fig. 3 presents one reconstructed image (using the same $x_{T}$) with three variations (using random $x_{T}$). The results show that the majority of semantic information in the images is accurately recreated, confirming that the prototypes effectively capture meaningful semantic features.
Figure 3: Reconstruction and variations with fixed $s$, random $x_{T}$.
Figure 4: Interpolation between two images. First row: CheXpert [12], from 75-year-old female $w/o$ enlarged heart (left) to 27-year-old male $w/$ enlarged heart (right). Last two rows: CelebA.
@@
Given two images $x_{0}^{1}$ and $x_{0}^{2}$, we first retrieve its corresponding prototype activation vector and starting noise by reverse DDIM process: $(s^{1},x_{T}^{1})$ and $(s^{2},x_{T}^{2})$, and then generate new samples using $(\mathrm{Lerp}(s^{1},s^{2};t),\mathrm{Slerp}(x_{T}^{1},x_{T}^{2};t))$ for steps $t\in[0,1]$, where $\mathrm{Lerp}$/$\mathrm{Slerp}$ represents linear/spherical linear interpolation respectively. Results are shown in Fig. 4.
Figure 5: Visualization of selected prototypes and their semantic interpretations. Here, $x_{0}$ denotes the original image, $\hat{x}_{0,s^{\prime}}$ denotes the generated image guided by condition $s^{\prime}$, where $s^{\prime}$ is the enhanced prototype activation vector on $j$-th prototype. Red rectangle highlights the most activated patch in $\hat{x}_{0,s^{\prime}}$, which is considered as the visual representation of the chosen prototype, also shown in the third row. Note that prototype semantics are not pre-annotated but inferred through observation.
Visualization of the learned prototypes.
A key distinction of our proposed method compared to other autoencoder-based diffusion models lies in the intrinsic interpretable semantic latent through visualization. Ideally, by training, each prototype inherently represents distinct content, and our approach is explicitly designed to allow visual interpretation of these representations. In practice, we maximize the similarity score of selected prototype and identify the image patch where each prototype responds most strongly – a process we refer to as extrapolation, in contrast to interpolation. Detailed steps are provided in Sec. 3.3. In Fig. 5 we visualize selected prototypes and listed their semantic meanings from observation on the top.
Consistency of prototype visualization.
Fig. 6 shows prototype visualization across four random samples for a fixed prototype. The most activated patches, representing the learned prototypes, remain consistent across samples.
Figure 6: Consistency of prototype visualization. We randomly select four samples and generate their enhanced images conditioned on prototype $j$. The most activated patch, representing the chosen prototype, remains consistent across all samples.
@@
By adjusting the condition $s$ we can edit the image with specific semantic requests. As shown in Fig. 1, we enhance features like “white collar” and “curly hair” and diminish “smile” by modifying the corresponding prototype similarity scores. Furthermore, pushing a selected dimension of $s$ to extreme values (ranging from 0.0 to 3.0) results in a smooth and continuous enhancement of the associated semantic information, as demonstrated in Fig.7. Unlike interpolation, which remains within the observed range, this process extends beyond the original data distribution, making it an extrapolation.
Figure 7: Extrapolation of prototype magnitude between 0 to 3.
@@ 4.2 Prototype Quality
Following [33], we test the prototype quality via a downstream classification task on the prototype activation vectors $s$ using a logistic regression classifier trained with 5-fold cross validation, reporting AUROC in Tab. 1 & 2. Our model outperform 3 out of 4 datasets in latent (prototype) quality, with particularly strong performance on CelebA and FFHQ. The lower latent quality for FMNIST may stem from Patronus prioritizing localized features, while DiffAE and InfoDiff emphasize global structures, which better capture the semantic information of FMNIST due to its high inter-class variability. For the datasets where semantic information is more localized, Patronus achieves a marked improvement.
@@
We quantify the disentanglement on CelebA using TAD [37]. Following [33], we first remove the highly correlated attributes, then compute the AUROC score for each dimension of the prototype activation vector $s$. An attribute is “captured” if any dimension achieves AUROC $>$ 0.75. TAD is the sum of AUROC differences between the top two predictive dimensions per captured attribute. As shown in Tab. 1, Patronus outperforms previous models in both TAD and captured attributes.
@@ 4.3 Generation Quality
FMNIST Cifar10 FFHQ
Latent FID $\downarrow$ Latent FID $\downarrow$ Latent FID $\downarrow$
@@
4.4 Diagnosing Generative Models with Prototypes
Figure 8: Patronus captures unwanted correlations of attributes (bias). We present four samples with their generated outputs, enhanced on a smile-related prototype and three hair-color prototypes. Prototypes are selected based on observation, with correctness confirmed by choosing the highest AUROC (bold value after ”/”) for single-dimension prediction. The Spearman correlation between hair-color and smile prototypes, shown in green (positive) and red (negative), highlights dataset bias. It is interesting to see that enhancing the black hair prototype shifts generated images from smiling to non-smiling, and vice versa. Noted that the reported values are computed on the entire test set.
We manipulated a subset of CelebA to introduce an unwanted correlation between hair color and smile: all blonde/brown-haired images smile, while black-haired ones do not. Fig. 8 visualizes the prototype for “smile” alongside one for each hair color categories. The parentheses list the Spearman correlation with the “smile” prototype, and the AUROC score for predicting the hair color using this prototype similarity score alone. Results confirm the model captures the introduced bias, as smile positively correlates with blonde/brown hair but not black. Prototypes for “smile” and hair colors were chosen based on the highest AUROC for single-dimension $s$ prediction.
Interestingly, enhancing a single prototype reveals unwanted correlations in the generated images. For example, increasing the black hair prototype shifts the smile property from “smile” to “non-smile” and vice versa (see Fig. 8). Similar patterns emerge without subset manipulation. In Fig. 7, increasing “white collar” enhances male features, while decreasing “eye makeup” reduces female features.
This emphasizes how Patronus can be utilized to discover unwanted model behavior.
@@ 5 Discussion and Conclusion
Given a generated image, Patronus reveals when each prototype emerges in the generation process by obtaining the prototype similarity score from estimated $\hat{x}_{0}$ at each timestep (more details in Appendix), as shown in the top-right corner of Fig. 1. Thus, by subtracting the two time-sequential $s$ of the semantically-enhanced image $\hat{x}_{0,s\prime}$ and $x_{0}$, we illustrate how each prototype emerges temporally in the diffusion process. As shown in Fig. 9, none of the prototypes have a significant emergence in the first 200 stages of generation. Interestingly, prototypes relating to lower spatial frequency attributes appear earlier during generation, such as “wearing red”; while higher spatial frequency attributes, like “curly hair”, emerge later in the generation process.
Figure 9: Different prototypes emerge at different stages of the diffusion process. The similarity score improvement between an edited image targeting prototype $j$ and the original image reflects how $p_{j}$ emerges over time. Here we show the average and standard deviation (as shadow) of similarity score improvement from 100 randomly generated samples. See text for a discussion on results.
@@
We test whether correlation between prototypes is caused by prototype collapse, where multiple prototypes represent the same semantics. To assess this, we introduce a Prototype Distinct Loss to encourage prototype disentanglement and evaluate its impact compared to using the denoiser loss alone. The Prototype Distinct Loss is defined as: $\mathcal{L}_{distinct}=\frac{1}{N}\sum_{i=1}^{N}\max(0,\delta-\min_{j\neq i}D_{ij})$, where $D$ is the cosine distance with absolute similarity: $D_{ij}=1-|\frac{p_{i}\cdot p_{j}}{\|p_{i}\|\|p_{j}\||}|$. The margin $\delta$ is set to 0.5 and 1.0. Notably, setting $\delta=1.0$ enforces prototypes to be orthogonal. We initialize the new model using the original network parameters and train it for an additional 100 epochs using $L_{ddpm}+L_{distinct}$. Fig. 10 shows that the new models do not see a substantial change in the learned prototypes, suggesting that the prototypes optimized via the denoising objective are already sufficiently decorrelated without explicit regularization.
(a) Distribution of Pearson correlation values for prototype similarity scores, computed from 1,024 randomly selected training samples. Results compare the original training and training with an additional loss term. The difference is minimal, with only a small fraction of pairs exceeding an absolute correlation of 0.5 (red dashed lines) for all cases.
(b) Prototype visualizations and their semantic interpretations remain consistent regardless of the additional loss.
Figure 10: Effect of additional loss on prototype diversity. (Top) Pearson correlation distribution of prototype similarity scores shows minimal differences regardless of additional loss. (Bottom) Prototype visualizations remain consistent, indicating negligible impact on semantic representations.
Limitations of Cross-Modal Interpretability.
We emphasize interpreting diffusion models without cross-modality by design for two key reasons:
  • Language’s limitation in representing complexity: While language, as a discursive symbolism, serves as a powerful medium for interpretation, it alone cannot fully represent non-symbolic sensory and semantics complexity [17]. This is evident in the superior performance of multi-modal learning over single-modality; prior work [6, 9] further shows the restricted perceptual capacity of language-only generation by introducing visual cues.
  • Inherited bias from the text embedding: (1) Incomplete text representations: if certain concepts are not explicitly named (e.g., medical devices in radiology reports), the model cannot learn their visual counterparts. (2) Spurious correlations: text data may encode unintended biases, such as differing report detail levels by patient demographics, which may propagate into generated images.

### v2

@@ 4 Experiments and Results
In our experiments, we first assess the semantic meaning of learned prototypes (Sec. 4.1) through reconstruction, interpolation, extrapolation, and manipulation tasks. We also investigate prototype visualization and their consistency. We then analyze prototype quality and generation performance (Sec. 4.24.3). Finally, we explore how Patronus aids in diagnosing shortcut learning in diffusion models by identifying potentially unwanted correlations via prototype similarity scores (Sec. 4.4).
We use five datasets: Quantitative evaluation on FMNIST (xiao2017fashion), CIFAR-10 (krizhevsky2009learning), FFHQ (karras2019style); qualitative analysis on CheXpert (irvin2019chexpert); and in-depth quantitative and qualitative experiments on CelebA (liu2015faceattributes). We use 100 prototypes for all datasets except FMNIST, which has 30. Prototypes are encoded with a shape of (1,1,128).
@@
We firstly verify that the learned prototypes are semantically meaningful by the following analyses.
Reconstructing image semantics from prototype activations.
We extract the prototype activation vector $s=Enc(x_{0})$ and sample random noise $x_{T}\sim\mathcal{N}(0,\mathbf{I})$ to generate new images $\hat{x}(s,x_{T})$. Fig. 2a presents one reconstructed image (using the same $x_{T}$) with three variations (using random $x_{T}$). The results show that the majority of semantic information in the images is accurately recreated, confirming that the prototypes effectively capture meaningful semantic features.
Figure 2: Assessing the semantic meaning of the learned prototypes. (a) Reconstruction. (b) Extrapolation. (c) Interpolation. Left: CheXpert, from 75-year-old female $w/o$ enlarged heart (left) to 27-year-old male $w/$ enlarged heart (right). Right: 2 examples from CelebA.
@@
Given two images $x_{0}^{1}$ and $x_{0}^{2}$, we first retrieve its corresponding prototype activation vector and starting noise by reverse DDIM process: $(s^{1},x_{T}^{1})$ and $(s^{2},x_{T}^{2})$, and then generate new samples using $(\mathrm{Lerp}(s^{1},s^{2};t),\mathrm{Slerp}(x_{T}^{1},x_{T}^{2};t))$ for steps $t\in[0,1]$, where $\mathrm{Lerp}$$\mathrm{Slerp}$ represents linear/spherical linear interpolation respectively. Results are shown in Fig. 2c.
Figure 3: Prototype visualization and consistency. (a) Prototype visualization. Here, $x_{0}$ denotes the original image, $\hat{x}_{0,s^{\prime}}$ denotes the generated image guided by condition $s^{\prime}$, where $s^{\prime}$ is the enhanced prototype activation on $j$-th prototype. Red square highlights the most activated patch, which is considered as the visual representation of the chosen prototype, also shown in the third row. Note that prototype semantics are not pre-annotated but inferred through observation. (b) Visualization across random samples demonstrates that each prototype encodes consistent semantics.
Visualization of the learned prototypes and consistency validation.
Unlike other autoencoder-based diffusion models, our approach is explicitly designed to yield an interpretable semantic latent space, where each prototype is trained to capture distinct semantic content. In practice, we realize this by amplifying a prototype’s activation and extracting the most responsive patch (see Sec. 3.3). Fig. 3a shows selected prototypes with their semantic visualizations, and Fig. 3b demonstrates that the most activated patches remain semantically consistent across samples.
@@
By adjusting the condition $s$ we can edit the image with specific semantic requests. Furthermore, pushing a selected dimension of $s$ to extreme values (ranging from 0.0 to 3.0) results in a smooth and continuous enhancement of the associated semantic information, as demonstrated in Fig.2b. Unlike interpolation, which remains within the observed range, this process extends beyond the original data distribution, making it an extrapolation.
@@ 4.2 Prototype Quality
We test the prototype quality via a downstream classification task on the prototype activation vectors $s$ using a logistic regression classifier trained with 5-fold cross validation, reporting AUROC in Tab. 1 & 2. Our model outperform 3 out of 4 datasets in latent (prototype) quality, with particularly strong performance on CelebA and FFHQ. The lower latent quality for FMNIST may stem from Patronus prioritizing localized features, while DiffAE and InfoDiff emphasize global structures, which better capture the semantic information of FMNIST due to its high inter-class variability. For the datasets where semantic information is more localized, Patronus achieves a marked improvement.
@@
We quantify the disentanglement on CelebA using TAD (yeats2022nashae). Following wang2023infodiffusion, we first remove the highly correlated attributes, then compute the AUROC score for each dimension of the prototype activation vector $s$. An attribute is “captured” if any dimension achieves AUROC >> 0.75. TAD is the sum of AUROC differences between the top two predictive dimensions per captured attribute. As shown in Tab. 1, Patronus outperforms previous models in both TAD and captured attributes.
@@ 4.3 Generation Quality
FMNIST CIFAR-10 FFHQ
Latent AUROC $\uparrow$ FID $\downarrow$ Latent AUROC $\uparrow$ FID $\downarrow$ Latent AUROC $\uparrow$ FID $\downarrow$
@@
4.4 Diagnosing Shortcut Learning in Diffusion Models
Figure 4: Patronus as an interpretable tool for (a) Detecting shortcut learning. Enhancing hair-color prototypes reveals their correlation with other attributes (e.g., smile), thereby exposing unwanted biases. (b) Tracing prototype emergence. Different prototypes appear at different stages of the diffusion process, as indicated by similarity score improvements, which provides insights for more effective image editing strategies.
We manipulated a subset of CelebA to introduce an unwanted correlation between hair color and smile: all blonde/brown-haired images smile, while black-haired ones do not. Results confirm the model captures the introduced bias, as increasing the black hair prototype shifts the smile property from “smile” to “non-smile” and vice versa (see Fig. 4a). Similar patterns emerge without subset manipulation. In Fig. 2b, decreasing “eye makeup” reduces female features. This emphasizes how Patronus can be utilized to discover unwanted model behavior, such as shortcut learning in this case.
@@ 5 Discussion and Conclusion
Given a generated image, Patronus reveals when each prototype emerges in the generation process by obtaining the prototype similarity score from estimated $\hat{x}_{0}$ at each timestep (more details in Appendix), as shown in the top-right corner of Fig. 1a. Thus, by subtracting the two time-sequential $s$ of the semantically-enhanced image $\hat{x}_{0,s\prime}$ and $x_{0}$, we illustrate how each prototype emerges temporally in the diffusion process. As shown in Fig. 4b, none of the prototypes have a significant emergence in the first 200 stages of generation. Interestingly, prototypes relating to lower spatial frequency attributes appear earlier during generation, such as “wearing red”; while higher spatial frequency attributes, like “curly hair”, emerge later in the generation process. This insight could potentially improve the efficiency of image editing or counterfactual generation by guiding how far an image should be reversed in the diffusion process. It could also support bias mitigation by leveraging the same mechanism once unwanted correlations are detected.
@@
[…]